# Object independent scatter sensitivities for PET, applied to scatter estimation through fast Monte Carlo simulation

### v1

@@ Abstract
Main result In scenario 1, scatter-compensated reconstructions achieved $1\%$ global bias in all active regions relative to true-only reconstructions. In scenario 2, noisy scatter estimates caused strong positive bias, but Gaussian smoothing restored accuracy to scenario 1 levels. In scenario 3, joint estimation under low-count conditions maintained $1\%$ global bias in nearly all regions.
@@ 1 Introduction
In most cases (barring energy-based [1, 2] and some deep learning-based [3] methods), the scatter estimate is calculated iteratively, by reconstructing an estimate of the true activity distribution with the current scatter estimate, simulating a new scatter estimate with this distribution, and then repeating this process a few times. In clinical practice, the scatter estimates are simulated using tail-fitted single-scatter simulation. This procedure makes use of a few shortcuts to ensure the computation can be done within a time frame that is acceptable for routine clinical use, while still producing good results in almost all cases [4, 5, 6]. Monte Carlo (MC) simulation is an alternative to tail-fitted single-scatter simulation that is believed to be more accurate and more robust to noise in the acquisition sinogram, due to the fact that it models most of the physical processes that take place during a scan. MC Simulators like Gate [7] and SimSet [8] can simulate very accurate scatter estimates [9], but require a lot of computation time to do so, and thus are not feasible for use in clinical practice. Recently though, very fast MC simulators were developed [10, 11, 12], and their feasibility for scatter estimation within clinically acceptable times was demonstrated [13].
@@ 2.1 Simulated scans in GATE
Instead of real PET acquisitions, we evaluated our approach on acquisitions using a replica of the GE Signa PET-MR scanner [16] built in GATE, in order to have control over all scanner parameters, and to have access to the true activity images and coincidences. In total, acquisitions were simulated for 3 different phantoms (Figure 2). In these simulated scans, we used 11.2% as the energy resolution, and 425-650 keV as the energy window. TOF resolution was set to 380 ps FWHM, similar to the real scanner. Since accurate methods for randoms estimation and compensation are available and widely used, randoms were not simulated for our evaluation, only trues and scatters. The effect of decay on activity concentration was not simulated.
@@ 2.2.2 Photon detection probability LUT
To evaluate the effects of each LUT component, lower dimensional LUTs were computed by summing over the unused dimensions of the actor and detector histograms (taking into account whether coincidences in gaps would be kept during listmode processing or not), and then dividing them as before. We created a 1D LUT using just energy information, one 3D LUT containing energy and angle information and another 3D LUT containing energy and crystal information. The processing of coincidences hitting gaps in the MCGPU-PET listmode would be handled slightly differently for each LUT. For LUTs without crystal information (or in the absence of a LUT), coincidences containing photons that hit a gap were discarded. Therefore, photons hitting gaps were discarded from the actor and detector histograms before creating those LUTs. For the 3D histogram with crystal information, photons hitting a gap would be assigned to the nearest edge crystal after receiving a probability. For the 5D LUT, this assignment was done by a second LUT.
@@ 2.2.3 Edge crystal assignment LUT
In summary, the listmode processing (with the LUT) takes the original MCGPU-PET listmode (using cylinder coordinates) as the input, and outputs a new listmode file that uses crystals instead of coordinates. This processing either assigns a weight to an event or discards it. It never converts a single input event to more than one weighted output events, so the number of events never increases. This is why gap events are not distributed over the edge crystals. During reconstruction or unlisting to a sinogram, the event contributes that weight, which is a value between 0 and 1.
@@ 3.1 Known activity distribution, unlimited simulation time
In the first scenario, we tested the combination of different LUTs with their respective scatter sensitivity sinograms for scatter estimation when the activity distribution is known, and when it is possible to use long, high count MC simulations to obtain high-SNR scatter estimates. Figure 8 shows cross-sections through the 5D detection probability LUT. The MCGPU-PET scatter simulations produced $6\cdot 10^{10}$ scatter coincidences in each simulation, which translates to around 36 hours of real-world simulation time per phantom and per LUT on our Nvidia A6000 GPUs. The intermediate and final scatter sinograms for the small cylinder are shown in Figure 9. This figure also shows a comparison with the scatter sinogram from the GATE simulation, and the respective scatter sensitivity sinograms for each LUT. The scatter estimate generated using the 5D LUT and scatter sensitivity shows the smallest differences with the GATE scatter sinogram.
[…]
[width=0.75]Figures/Probabilities.png
@@
[…]
[width=1.0]Figures/Recons.png
Figure 10: Differences between reconstructions of the small cylinder phantom made using MC-simulated scatter sinograms (with different LUTs and scatter sensitivities), the trues only reconstruction (TO) and the ground truth image (GT). The differences are expressed as a percentage of the phantom concentration (background concentration for the NEMA phantom). Recons are averaged over all planes for the cylindrical phantoms, a single slice is shown for the NEMA phantom. The differences are expressed as a percentage of the phantom concentration. Note the different scales.
@@
Table 1: MSE and bias of the activity values compared to the trues only reconstruction per ROI, for each phantom and LUT combination. (CL: Cylinder, BS: Biggest sphere, BG: Background, CA: Cavity)
@@ 3.2 Known activity distribution, very limited simulation time
For the second scenario, we first investigated the effect of the number of MC simulated counts on the scatter estimates by performing reconstructions with scatter estimates of different count levels. The scatter-compensated reconstructions in Figure 11 show that lower count levels (lower SNR) in the MCGPU-PET simulations increases the bias in the resulting reconstructions.
[…]
[width=]Figures/snr_recons.png
@@ 4.1 Detector efficiency LUT
The 5D LUT obtained from the GATE simulations was still slightly noisy, though this did not have a noticeable effect on the sinograms it produced. The LUT also captured some interesting effects, which can be seen in the 2D slices in Figure 8:
@@ 4.2 Known activity distribution, unlimited simulation time
The comparison with the GT is only shown in Figures 10 and 11 and left out of the ROI comparison, as we believe it is of little added value. The reason is that our trues only reconstruction contains a slight negative bias $0.5\%$ compared to the GT activity image, likely due to an attenuation modeling mismatch between our forward model and the GATE simulation. We found a similar bias when performing a scatter-compensated reconstruction both using the exact GATE scatter sinogram and when using a higher SNR scatter sinogram with 14 radial symmetries applied. Therefore, we believe it is more logical to compare the scenario 1 reconstructions to the trues only reconstruction, as the other reconstructions will contain a similar bias caused by this forward model mismatch.
The leftover positive bias in the 5D LUT reconstruction has two possible causes related to the scatter sensitivity sinogram. One possible explanation is the fact that the scatter sensitivity sinogram is also slightly noisy, because it is calculated from a GATE scatter sinogram. This noise then propagates into the scatter estimates, and as demonstrated in Figure 11, even very small amounts of noise can cause bias in the reconstructions. However, running the simulation that produces the GATE scatter sinogram for longer to get a better SNR was not feasible, as the runtime for such high-SNR GATE simulations quickly reaches weeks or months. Another possibility is that the LUT is missing an object-dependent effect which in turn creates bias in the scatter sensitivity sinogram.
Indeed, it is not realistic to expect that a LUT in combination with a sensitivity correction sinogram can perfectly model the physics that are simulated in advanced MC simulators like GATE or SimSet. However, our method for incorporating the scatter sensitivities is a valid approximation for simulators which model a simple cylindrical scanner and do not simulate detector physics (i.e. detect all events hitting the cylinder). It can even be extended to different scanner geometries, which would only require repeating the simulations for the LUT and trues and scatter sensitivity sinograms with the new geometry. That being said, the value of a fast simulator which can simulate detector physics [12] should not be underestimated.
@@ 4.4 Unknown activity distribution, very limited simulation time
The smoothing approach and the included normalization were purposefully kept simple as well, for the same reasons as the iterative scatter estimation. The “smoothing normalization sinogram” is not entirely object independent, and leaves some high-frequent structures in the sinograms after division. Smoothing them away causes reconstruction artifacts to appear, as demonstrated in Figure 12. More sophisticated normalization and/or smoothing methods may be able to avoid smoothing away these structures.
@@ 4.5 Future work
The logical next step after the development of this scatter normalization and estimation approach would be to apply it to scans of a real PET scanner (e.g. the modeled Signa PET/MR). We believe there are two main roadblocks before our approach to scatter estimation can be used on real scanners:
  1. 1.
    When working with real scans, aside from randoms, one must also account for effects like dead time, pile-up corrections and relative crystal sensitivities (which may be different for trues and scatters), some of which are accounted for in vendor software using proprietary algorithms.
  2. 2.
    Scanner features like the recovery of inter-crystal scatter between detector blocks [23, 24] complicate the calculation of the LUT, as this type of advanced coincidence processing is not natively available in GATE.
As stated before, these extra variables, coupled with the lack of control over the inputs to the scatter estimation pipeline, are the reason that we have not included a section on real scans in this paper. Now that we have achieved proof-of-concept, we will investigate how to include these extra variables in our scatter estimation pipeline.
Once we have successfully applied our scatter estimation approach to a real scan, we will also further optimize the speed of our iterative scatter estimation pipeline, to make it more competitive with other scatter estimation methods.

### v2

@@ Abstract
We also adapted the approach to test it on two acquisitions from a real Signa PET/MR.
Main result In scenario 1, scatter-compensated reconstructions achieved $1\%$ global bias in all active regions relative to true-only reconstructions. In scenario 2, noisy scatter estimates caused strong positive bias, but Gaussian smoothing restored accuracy to scenario 1 levels. In scenario 3, joint estimation under low-count conditions maintained $1\%$ global bias in nearly all regions. For real scans, the Monte Carlo-based scatter estimate was very similar to the vendor scatter estimate.
@@ 1 Introduction
In most cases (barring energy-based [8, 6] and some deep learning-based [21] methods), the scatter estimate is calculated iteratively, by reconstructing an estimate of the true activity distribution with the current scatter estimate, simulating a new scatter estimate with this distribution, and then repeating this process a few times. In clinical practice, the scatter estimates are simulated using tail-fitted single scatter simulation (SSS). This procedure makes use of a few shortcuts to ensure the computation can be done within a time frame that is acceptable for routine clinical use, while still producing good results in almost all cases [20, 27, 22]. Monte Carlo (MC) simulation is an alternative to tail-fitted SSS that is believed to be more accurate and more robust to noise in the acquisition sinogram, due to the fact that it models most of the physical processes that take place during a scan. MC Simulators like Gate [14] and SimSet [9] can simulate very accurate scatter estimates [3], but require a lot of computation time to do so, and thus are not feasible for use in clinical practice. Recently though, very fast MC simulators were developed [10, 18, 23, 16, 7, 5], and their feasibility for scatter estimation within clinically acceptable times was demonstrated [19].
@@
  5. 5.
    Given some additional modifications to the simulator, the same fast, low-count MC simulations can be used to estimate scatter for real scans, without the need for tail fitting.
@@
After the quantitative evaluation on simulated acquisitions, we included the effect of inter-crystal scatter recovery, a correction factor for non-ideal detectors and the dead-time and pile-up corrections in the scatter estimation pipeline. We then estimated the scatter for two phantom acquisitions on a real Signa PET/MR and compared the resulting sinograms and reconstructions to those obtained using tail-fitted SSS from the vendor.
@@ 2.1 Simulated scans in GATE
We first evaluated our approach on acquisitions using a replica of the GE Signa PET-MR scanner [17] built in GATE, in order to have control over all scanner parameters, and to have access to the true activity images and coincidences. In total, acquisitions were simulated for 3 different phantoms (Figure 2). In these simulated scans, we used 11.2% as the energy resolution, and 425-650 keV as the energy window. TOF resolution was set to 380 ps FWHM, similar to the real scanner. Since accurate methods for randoms estimation and compensation are available and widely used, randoms were not simulated for our evaluation, only trues and scatters. The effect of decay on activity concentration was not simulated.
@@ 2.2.2 Photon detection probability LUT
To evaluate the effects of each LUT component, lower dimensional LUTs were computed by summing over the unused dimensions of the actor and detector histograms (taking into account whether coincidences in gaps would be kept during listmode processing or not), and then dividing them as before. We created a 1D LUT using just energy information, one 3D LUT containing energy and angle information and another 3D LUT containing energy and crystal information. The processing of coincidences hitting gaps in the MCGPU-PET listmode would be handled slightly differently for each LUT. For LUTs without crystal information (or in the absence of a LUT), coincidences containing photons that hit a gap were discarded. Therefore, photons hitting gaps were discarded from the actor and detector histograms before creating those LUTs. For the 3D LUT with crystal information, photons hitting a gap would be assigned to the nearest edge crystal after receiving a probability. For the 5D LUT, this assignment was done by a second LUT.
@@ 2.2.3 Edge crystal assignment LUT
In summary, the listmode processing (with the LUT) takes the original MCGPU-PET listmode (using cylinder coordinates) as the input, and outputs a new listmode file that uses crystals instead of coordinates. This processing either assigns a weight to an event or discards it. It never converts a single input event to more than one weighted output event, so the number of events never increases. This is why gap events are not distributed over the edge crystals. During reconstruction or unlisting to a sinogram, the event contributes that weight, which is a value between 0 and 1.
@@ 2.3.4 Evaluation metrics
2.4 Scatter estimation for phantom scans on the GE Signa PET/MR
After the quantitative evaluation on the simulated acquisitions, we also applied our scatter estimation approach to real acquisitions, as a proof-of-concept. For this purpose, we acquired two scans for phantoms similar to the ones in the simulated evaluation. One big difference between these acquisitions and the simulated ones was the presence of random coincidences, for which a separate expectation sinogram was estimated using vendor-provided software. The scanner geometry, energy resolution, energy window and TOF resolution were identical to the simulated acquisitions.
2.4.1 Small cylinder acquisition
The first phantom acquisition consisted of a small cylinder with a length of 20 cm and a radius of 10 cm, placed in the center of the scanner’s FOV. It was filled with 0.9 kBq/ml of F-18 at the start of the scan. Total scan time was 2 hours, resulting in $3.74\cdot 10^{8}$ prompt coincidences.
2.4.2 NEMA phantom acquisition
The second phantom acquisition consisted of a NEMA image quality phantom, placed in the center of the scanner’s FOV. The smallest and largest spheres were filled with an activity ratio of 4:1 compared to the background. The exact F-18 activity concentration was unfortunately not known, but estimated to be approximately 2.5 kBq/ml for the background and 10 kBq/ml for the spheres. Total scan time was 1 hour, resulting in $1.22\cdot 10^{9}$ prompt coincidences.
2.4.3 Scatter estimation for real acquisitions
To estimate scatter for real acquisitions, we had to account for several effects/phenomena present in our scanner, but not yet modeled during our simulated evaluation:
  1. 1.
    Decay: The effect of activity decay was not directly applied during the MCGPU-PET simulation, but rather incorporated as a single scale factor based on the length of the scan, for practical reasons.
  2. 2.
    Random coincidences: The randoms expectation sinogram was estimated using vendor software, and included in the forward model during reconstruction.
  3. 3.
    Inter-crystal scatter recovery: In the real scanner, photons that scatter inside the detector crystals between neighboring blocks are recovered [12, 25], an effect we did not simulate in our GATE acquisitions. The resulting sensitivity increase for the scanner was incorporated in the LUT by including the recovered photons (in GATE) in the probability calculations, and the corresponding scatter sensitivity sinogram was obtained by re-running the required simulations with the updated LUT and GATE logic.
  4. 4.
    Dead-time and pile-up: The vendor estimates for dead-time and pile-up were included in the simulator’s forward model.
  5. 5.
    Different relative detector sensitivities: The vendor estimates for non-uniformity of the detector sensitivities were included in the simulator’s forward model.
  6. 6.
    Correction factor for non-ideal detectors: In our testing, even with all the modifications above, the ratio between the expected number of scattered coincidences for both phantom acquisitions and the number we simulated was only around %75. We assume this is due to the fact that the detectors in GATE (which also determine the detection probabilities in the LUT) are an ideal version of those in the scanner, and that this difference should be captured by the trues sensitivity sinogram of the scanner. Since the sensitivity of the scanner is not constant over time, we estimated one global scale factor per acquisition from the ratio of the GATE trues sensitivity sinogram and the (acquisition-specific) scanner trues sensitivity sinogram, which assumes that this global scale factor is the same for true and scattered coincidences. The GATE trues sensitivity sinogram used for this purpose was different from the one used for reconstructing our simulated acquisitions, as it includes corrections for the effects mentioned above (similar to the scanner trues sensitivity sinogram). The resulting correction factors were 0.786 for the cylinder and 0.730 for the NEMA phantom, respectively.
Apart from these modifications, the scatter simulation and joint estimation approach was identical to the one used in the third scenario of the simulated evaluation (unkown activity, short simulation time).
2.4.4 Comparison to tail-fitted SSS
Aiming for the fairest possible comparison between tail-fitted SSS and MC simulated scatter, we kept as much as possible of the reconstruction pipeline identical between both approaches, and only replaced the scatter estimate. The randoms estimate, the trues sensitivity sinogram, the attenuation sinogram etc. were the same for both approaches. The reconstruction parameters were identical to those in the simulated evaluation (i.e. TOF-OSEM, 4 iterations and 28 subsets, $192\times 192\times 90$ voxels of $2.734\times 2.734\times 2.890\ \mathrm{mm^{3}}$). Since the Signa is a PET/MR scanner, attenuation maps for both phantoms had to be obtained from CT-derived attenuation templates, and had to be combined with an attenuation template of the scanner bed. The estimated scatter sinograms for both approaches were compared by plotting non-TOF radial profiles.
@@ 3.1 Known activity distribution, unlimited simulation time
In the first scenario, we tested the combination of different LUTs with their respective scatter sensitivity sinograms for scatter estimation when the activity distribution is known, and when it is possible to use long, high count MC simulations to obtain high-SNR scatter estimates. Figure 8 shows cross-sections through the 5D detection probability LUT. The MCGPU-PET scatter simulations produced $6\cdot 10^{10}$ scatter coincidences in each simulation, which translates to around 36 hours of real-world simulation time per phantom and per LUT on our Nvidia A6000 GPUs. The intermediate and final scatter sinograms for the small cylinder are shown in Fig. 9. This figure also shows a comparison with the scatter sinogram from the GATE simulation, and the respective scatter sensitivity sinograms for each LUT. The scatter estimate generated using the 5D LUT and scatter sensitivity shows the smallest differences with the GATE scatter sinogram.
@@
Figure 10: Differences between reconstructions of all phantoms made using MC-simulated scatter sinograms (with different LUTs and scatter sensitivities), the trues only reconstruction (TO) and the ground truth image (GT). The differences are expressed as a percentage of the phantom concentration (background concentration for the NEMA phantom). Recons are averaged over all planes for the cylindrical phantoms, a single slice is shown for the NEMA phantom. The differences are expressed as a percentage of the phantom concentration. Note the different scales.
@@
Table 1: MSE and bias of the activity values compared to the trues only reconstruction per ROI, for each phantom and LUT combination. Values in bold indicate the lowest MSE and smallest bias for each region. (CL: Cylinder, BS: Biggest sphere, BG: Background, CA: Cavity)
@@ 3.2 Known activity distribution, very limited simulation time
For the second scenario, we first investigated the effect of the number of MC simulated counts on the scatter estimates by performing reconstructions with scatter estimates of different count levels. The scatter-compensated reconstructions in Fig. 11 show that lower count levels (lower SNR) in the MCGPU-PET simulations increases the bias in the resulting reconstructions.
@@ 3.3 Unknown activity distribution, very limited simulation time
3.4 Real scans on the Signa PET/MR
The iterative scatter estimation for the real scans was run for 10 iterations, using the same parameters for smoothing and the number of simulated counts as scenarios 2 and 3 in the simulated evaluation. The scatter estimation, including the final reconstruction, took around 3 hours. Of those 3 hours, 5 minutes were spent on MC simulation and 20 minutes on smoothing of the scatter sinograms. Reconstructions using the vendor and MC-simulated scatter, as well as sinogram profiles for the prompts (true, scattered and random coincidences) and contamination (scattered and random coincidences) sinograms are shown in Fig. 14. There are no apparent differences between the MC and vendor scatter reconstructions, and both contamination sinogram profiles also agree very well in most radial bins.
(a) Cylindrical phantom.
(b) NEMA phantom.
Figure 14: Plots of radial profiles through non-TOF sinograms averaged over all planes and 2 views, for the prompts and contamination sinograms. Large red arrows indicate mismatches between the prompts profile tail and both contamination profiles. Reconstructions made using the vendor scatter (SSS) and MC-simulated scatter are shown at the bottom. Small blue arrows show the orientation of the LORs contained in the respective views.
@@ 4.1 Detector efficiency LUT
The 5D LUT obtained from the GATE simulations was still slightly noisy, though this did not have a noticeable effect on the sinograms it produced. The LUT also captured some interesting effects, which can be seen in the 2D slices in Fig. 8:
@@ 4.2 Known activity distribution, unlimited simulation time
The comparison with the GT is only shown in Fig. 10 and 11 and left out of the ROI comparison, as we believe it is of little added value. The reason is that our trues only reconstruction contains a slight negative bias $0.5\%$ compared to the GT activity image, likely due to an attenuation modeling mismatch between our forward model and the GATE simulation. We found a similar bias when performing a scatter-compensated reconstruction both using the exact GATE scatter sinogram and when using a higher SNR scatter sinogram with 14 radial symmetries applied. Therefore, we believe it is more logical to compare the scenario 1 reconstructions to the trues only reconstruction, as the other reconstructions will contain a similar bias caused by this forward model mismatch.
The leftover positive bias in the 5D LUT reconstruction has two possible causes related to the scatter sensitivity sinogram. One possible explanation is the fact that the scatter sensitivity sinogram is also slightly noisy, because it is calculated from a GATE scatter sinogram. This noise then propagates into the scatter estimates, and as demonstrated in Fig. 11, even very small amounts of noise can cause bias in the reconstructions. However, running the simulation that produces the GATE scatter sinogram for longer to get a better SNR was not feasible, as the runtime for such high-SNR GATE simulations quickly reaches weeks or months. Another possibility is that the LUT is missing an object-dependent effect which in turn creates bias in the scatter sensitivity sinogram.
Indeed, it is not realistic to expect that a LUT in combination with a sensitivity correction sinogram can perfectly model the physics that are simulated in advanced MC simulators like GATE or SimSet. However, our method for incorporating the scatter sensitivities is a valid approximation for simulators which model a simple cylindrical scanner and do not simulate detector physics (i.e. detect all events hitting the cylinder). It can even be extended to different scanner geometries, which would only require repeating the simulations for the LUT and trues and scatter sensitivity sinograms with the new geometry. While our approach can provide a significant speedup of MC simulators by skipping the simulation of photon-detector interactions, the value of a fast simulator which can simulate detector physics should not be underestimated. We believe that once the speed of simulators with detector physics reaches a certain point, their superior accuracy will be preferable over the shorter simulation time of the detection probability LUT and sensitivity sinogram approach.
@@ 4.4 Unknown activity distribution, very limited simulation time
The smoothing approach and the included normalization were purposefully kept simple as well, for the same reasons as the iterative scatter estimation. The “smoothing normalization sinogram” is not entirely object independent, and leaves some high-frequent structures in the sinograms after multiplication. Smoothing them away causes reconstruction artifacts to appear, as demonstrated in Fig. 12. More sophisticated normalization and/or smoothing methods may be able to avoid smoothing away these structures.
@@
4.5 Real scans on the Signa PET/MR
Based on the reconstruction slices shown in Fig. 14, it is very hard to tell the difference between the reconstruction made using the vendor scatter, and the one made using MC-simulated scatter. In the sinogram profiles, we can see that the MC-based scatter sinogram follows the tails of the prompts sinogram (where only scatter is expected) very well. However, for the bins indicated with the red arrows, both contamination profiles fail to follow the prompt tails. These bins correspond to LORs passing through the bed of the scanner, so the mismatch might be explained by an imperfect estimate of the attenuation, which is plausible when using template-based attenuation images. At certain positions in the tails, the MC scatter performs better than the vendor scatter, which sometimes slightly overestimates the number of scatters. However, while the intrinsic scaling of the MC-based scatter estimation gives it a definitive advantage over tail-fitted SSS, it is impossible to make any statements about which is better (and under which conditions) without a proper reference and investigation. The results presented here merely serve to prove that it is possible to estimate scatter for real scans using MC simulation in a reasonably short time, and that the resulting sinograms and reconstructions are close to those produced by tail-fitted SSS (i.e. the current standard for clinical practice).
@@ 4.6 Future work
The scatter estimation pipeline we have presented here is a proof-of-concept, and there are a few aspects that future work could improve upon. One important point is the speed of the iterative scatter estimation approach and the processing of the MC simulation outputs, which we purposefully did not optimize, as we only aimed to show that scatter estimation can still be done accurately with fast, low-count MC simulations. Another point of improvement is the noise reduction step in the pipeline, as the current approach is fairly simple and may introduce additional artifacts. For real scans, it would also be useful to further investigate the reason for the differences between the GATE and scanner trues sensitivities, and to determine whether a single global scale factor is a reasonable approximation.
While out of the scope of this work, a deeper dive into the accuracy, robustness and speed of MC-based scatter estimation (as compared to tail-fitted SSS) will be essential if MC-based scatter estimation is ever to become the new clinical standard.